Solve linear equation using any technique.Name technique.check.
3x/4+6=-12
I would greatly appreciate your time and help.

Answer by rfer(16322)

associative property
3x/4+6=-12
3x/4+6-6=-12-6
3x/4=-18
3x/4*4=-18*4
3x=-72
x=-24
----------------
3(-24)/4+6=-12
-72/4+6=-12
-18+6=-12
-12=-12